Bvellla:

The number of buses needed for a school field trip depends on the total number of people going. One field trip required 7 buses for a total of 217 students and 35 adults. What is the rate of people per bus?

Bvellla:

PLEASE HELP

Aphmau:

217+35=252252/7=3636 people per bus